A chemist needs 80milliliters of a 25% solution but only has a 22% and a 30% solution available. find how many Milliliters of each solution should be mixed to get the desired solution
Answer by checkley77(12844)

.30x+.22(80-x)=.25*80
.30x+17.6-.22x=20
.08x=20-17.6
.08x=2.4
x=2.4/.08
x=30 ml. of 30% solution is used.
80-30=50 ml. of 22% solution is used.
Proof:
.30*30+.22*50=.25*80
9+11=20
20=20